OpenStudy (chupacabraj):

Find the exact area of the surface obtained by rotating the curve about the x-axis. 9x = y^2 + 27, 3 ≤ x ≤ 7

OpenStudy (chupacabraj):

Should I get y in terms of x first?

OpenStudy (phi):

I would not bother. notice we can write this as x= (1/9) y^2 + 3 3 <= x < 7 which is a ( shaped parabola if you like you could switch all the x and y's: y = (1/9) x^2 + 3 3 <= y <= 7 (corresponding to x=0 to 6) rotated about the y-axis

OpenStudy (chupacabraj):

using y = (1/9) x^2 + 3 3 <= y <= 7 I get \[SA = 2\pi \int\limits_{0}^{6}(\frac{ 1 }{ 9 }x^2+3)\sqrt{\frac{ 4 }{ 81 }x^2 +1} dx\]

OpenStudy (chupacabraj):

is this correct?

OpenStudy (phi):

using the "switched x-y version"

OpenStudy (phi):